OpenStudy (anonymous):

The sum of the measures of the remote interior angles of a triangle is half the measure of the angle adjacent to the exterior angle. find the measure of the exterior angle.

OpenStudy (anonymous):

i'm not sure, but i think the answer is 60 degrees.

OpenStudy (anonymous):

yes it actually is. how did you get it? :)

OpenStudy (anonymous):

let remote angles = x and y x + y = e 2(x+y) = 180-e dividing 2 = 180 -e / e solve this gives e = 60

OpenStudy (anonymous):

let us assume the angle adjacent to the exterior measures x, therefore, the sum of the remote interior angles will be 1/2x according to the property,"the sum of the measuresof the angles of a triangle is always 180", we get the following equation: 1/2x + X=180 2x + 1x/2=180 x= 180 * 2/3 x=120 but according to the equation, x is the angle adjacent to the exterior & it forms a linear pair with it. so, x + exterior angle =180 120 + exterior angle= 180 exterior angle=60